OpenStudy (anonymous):

how do I solve m=8/9,(2,-1)

OpenStudy (aaronq):

solve what exactly?

OpenStudy (anonymous):

the full question Find an equation of the line having the given slope and containing the given point. m=8/9,(2,-1) I don't know how to type the fraction 8/9 but that's what it is.

OpenStudy (aaronq):

okay, you know that m=slope, right? and the format for the point given is (x,y) so just plug those into y=mx+b

OpenStudy (aaronq):

and find b

OpenStudy (anonymous):

ok this is how I was doing it. Idk if I am setting this up right. If so this is where I get stuck. y-(-1)=(8/9) (x-2) y+1=8/9x 2/1= 16/9 this is where I get stuck how do I proceed

OpenStudy (aaronq):

you're not plugging the values in correctly, it's y=m*x+b m=8/9, (2,-1) (x,y) -1=(8/9)(2)+b now just solve for b

OpenStudy (anonymous):

thank you

OpenStudy (aaronq):

no problem !